OpenStudy (anonymous):

A box is to be constructed from a sheet of cardboard that is 10 cm by 50 cm by cutting out squares of length x by x from each corner and bending up the sides. What is the maximum volume this box could have? (Round your answer to two decimal places.)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

i already did it but i got it wrong 2 times please help

OpenStudy (phi):

is this calculus?

OpenStudy (anonymous):

yes

OpenStudy (phi):

did you draw a picture of what is going on ?

OpenStudy (anonymous):

u want me to show u what i have gotten so far?

OpenStudy (anonymous):

yes i did i got the equation (50-2x)(10-2x)(x)

OpenStudy (anonymous):

then all together 4x^3-120x^2+500x

OpenStudy (phi):

ok, now the derivative wrt x

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

then i found the maximum of the graph of 4x^3-120x^2+500x

OpenStudy (anonymous):

and i got 3.4978761

OpenStudy (phi):

you have \[ V = 4x^3-120x^2+500x \] and people find the derivative of that \[ \frac{dV}{dt} = 12x^2 -240x + 500 \] and set = to zero \[ 12x^2 -240x + 500=0\\ 3x^2 -60x+125=0 \]

OpenStudy (anonymous):

and now do i find the maximum of that?

OpenStudy (phi):

you find what x's make the equation 0 use the quadratic formula

OpenStudy (anonymous):

so now i use the quadratic formula or do i put it in the graphing calculator and find the maximum?

OpenStudy (phi):

we don't want the maximum of the derivative, we want to find the x value that makes the derivative 0. so we solve for x in 3x^2-60x+125=0 using the quadratic formula

OpenStudy (anonymous):

and then my answer from that is the final answer?

OpenStudy (anonymous):

i got this

OpenStudy (anonymous):

\[60+-\sqrt{-5100} all divided by 6

OpenStudy (phi):

you should get a real number

OpenStudy (anonymous):

i dont know how to simplify sqrt

OpenStudy (anonymous):

im really bad at it

OpenStudy (phi):

the number inside the square root should be -60*-60 - 4*3*125 that is not -5100

OpenStudy (anonymous):

i got 2100 sorry

OpenStudy (phi):

I would use a calculator and get a decimal answer. they say (Round your answer to two decimal places.)

OpenStudy (anonymous):

i am using a caluclator

OpenStudy (anonymous):

i did 4(3)125 and got 1500 then -60*-60 is 3600 and those 2 subtracted is 2100 right???

OpenStudy (phi):

yes so \[ x = \frac{60\pm\sqrt{2100}}{6} \]

OpenStudy (anonymous):

i got 60+10 sqrt{21} all divided by 6 is that right

OpenStudy (anonymous):

@phi

OpenStudy (phi):

that root will give you 10+ (sqrt(21))/6 which is bigger than 10 one of your sides is length 10, so x has to be smaller than 10 I would use the other root.

OpenStudy (anonymous):

i got 2.362373842

OpenStudy (phi):

yes, that looks ok. round to 2 decimal places.

OpenStudy (anonymous):

so 2.36 is the final answer

OpenStudy (phi):

yes